Sum of present ages of a man and his son is 60 years. Five years ago the mans age was 4 times his sons age. Then find their present ages.
Answer by mananth(16946)

Sum of present ages of a man and his son is 60 years. Five years ago the mans age was 4 times his sons age. Then find their present ages.
Let the son's age be x
Father's age will be 60-x
5 years ago
(60-x) -5 = (x-5) *4
300-5x -5= 4x-20
300+20-5=4x+5x
315= 9x
x=25 years
FAther's age = 60-25 = 35.